Two fishermen go ice fishing one morning. By noon, fisherman Dan has 6 fish and is catching 2 fish per hour. Fisherman Fred has 4 fish and is catching 3 per hour. Assume they will continue catching fish at the same rate.

A. Write a system of equations that relates the time to the total fish caught.

B. Solve your system using either substitution or elimination. Make sure to show all of your work for full credit.

C. Based upon your work in Part B, at what time will Fred and Dan have the same number of fish, and how many fish will they each have?

Answer by stanbon(75887)

Two fishermen go ice fishing one morning. By noon, fisherman Dan has 6 fish and is catching 2 fish per hour. Fisherman Fred has 4 fish and is catching 3 per hour. Assume they will continue catching fish at the same rate.
A. Write a system of equations that relates the time to the total fish caught.
Dan:: f(x) = 6 + 2x
Fred:: g(x) = 4 + 3x
------------------------------
B. Solve your system using either substitution or elimination. Make sure to show all of your work for full credit.
Note: You can only solve it at the time they have the same number of fish.
6+2x = 4+3x
x = 2 hrs
---
After 2 hrs f(x) = g(x) = 10 fish
----------------------------------
C. Based upon your work in Part B, at what time will Fred and Dan have the same number of fish,
Ans: 2 hrs after noot = 2:00 PM
------------------
and how many fish will they each have?
Ans: 10 fish
----------------
Cheers,
Stan H.
-----------